OpenStudy (anonymous):

Which pair of ratios forms a proportion? 10 over 16. and 5 over 32. 10 over 22. and 5 over 11. 14 over 8. and 7 over 16. 8 over 44. and 16 over 22.

OpenStudy (anonymous):

For there to be a proportion, if you place the fractions together and simplify each side, they should be equal, for example: 1) 10/16 = 5/32? 5/8 =/=5/32 2)10/22 = 5/11? 5/11=5/11 which is true so that's the proportion If that was too fast or something and you have a question about it, just ask